OpenStudy (anonymous):

Simpfily (2√5 + 3√7)^2 . Show your work. Justify each step. Need Help with Algebra I 9B Please!

OpenStudy (anonymous):

Justin Bieber needs help with math? does he even have to do math?

OpenStudy (anonymous):

anyways, you have to use the binomial expansion \[(2\sqrt{5}+3\sqrt{7})^2=(2\sqrt{5}+3\sqrt{7})((2\sqrt{5}+3\sqrt{7})\\ \quad=(2\sqrt{5})(2\sqrt{5})+(2\sqrt{5})(3\sqrt{7})+(3\sqrt{7})(2\sqrt{5})+(3\sqrt{7})(3\sqrt{7})\\ \quad=(4\times5)+(6\sqrt{35})+(6\sqrt{35})+(9\times7)\\ \quad=20+12\sqrt{35}+63\\ \quad=83+12\sqrt{35} \]

OpenStudy (anonymous):

(2√5 + 3√7)^2 \[(2\sqrt{5} )^{2}+(3\sqrt{7})^{2}+( 2\times2\sqrt{5}\times3\sqrt{7})\] \[4*5 +3*7 + 12\sqrt{35}\] \[20 +63+ 12\sqrt{35}\] \[83 + 12\sqrt{35}\]

OpenStudy (anonymous):

Disclaimer: the first post was for a quick laugh, guyz. So, laugh.

OpenStudy (anonymous):

Lol:)

OpenStudy (anonymous):

(2√5 + 3√7)^2=2^2*sqrt(5)^2+2*2 sqrt(5)*3^2*sqrt(7)+3^2*sqrt(7)^2=20+36*sqrt(35)+63=83+36*sqrt(35). Elaborated that time!!!

OpenStudy (anonymous):

Thank You Everyone i wish i could give everyone a medal...